# Towards a Quantum Information Theory of Hadronization: Dihadron Fragmentation and Neutral Polarization in Heavy Baryons

### v1

@@ VI TMD heavy-quark fragmentation
It is experimentally possible Brandenburg et al. (1997) that inclusive fragmentation $Q\to\Sigma_{Q}^{*}X$ in fact produces an isotropic spin state with $w_{1}\approx 2/3$. It is interesting to ask, therefore, whether breaking the azimuthal symmetry by measuring a transverse momentum $\bm{k}_{T}\perp\bm{z}$ provides access to richer spin-$1$ dynamics. Here the transverse momentum-dependent (TMD) observable $\bm{k}_{T}\gtrsim\Lambda_{\mathrm{QCD}}$, may e.g. be (a) the recoil of the heavy hadron pair in $e^{+}e^{-}\to\Sigma_{Q}^{*}H_{\bar{Q}}X$ von Kuk et al. (2023); Dai et al. (2024); von Kuk et al. (2024); Lee et al. (2025), (b) the hadron transverse momentum relative to a jet containing it Kang et al. (2020); Arratia et al. (2020); Lee et al. (2025), or (c) the transverse momentum of an additional fragmentation pion Bacchetta et al. (2023); Lee et al. (2025). In any of these cases, the most general light angular momentum state compatible with the symmetries reads Lee et al. (2025)
@@
where $k_{T}=\lvert\bm{k}_{T}\rvert$, $\bm{x}=\bm{k}_{T}/k_{T}$, $\bm{y}=\bm{z}\times\bm{x}$, and the $w^{(N,n)}$ are the nonperturbative coefficients.
Our previous discussion elegantly carries over if we define
@@
i.e., by finding the axis $\bm{z^{\prime}}$ that maximizes the neutral polarization rate $1-w_{1}^{\prime}$. A rotation from $\bm{z}$ to $\bm{z}^{\prime}$ in fact diagonalizes $\rho$ with eigenvectors $\lvert m_{z^{\prime}}\rangle$, so we can think of $\bm{z^{\prime}}$ as a novel polarization axis that is spontaneously generated by the fragmentation dynamics. Here we can make immediate use of $\bm{z}^{\prime}$ by repeating our previous QI analysis for $w_{1}^{\prime}$, as shown in Fig. 3, where $w^{(2,0)}=0$ corresponds to $w_{1}=2/3$ after integrating over $\bm{k}_{T}$. The value of $w_{1}^{\prime}$ is driven by $w^{(2,1)}$ and $w^{(2,2)}$, and even for $w_{1}=2/3$ we find physical regions in their parameter space with nontrivial QI interpretations $w_{1}^{\prime}<1/2$ or $w_{1}^{\prime}<w_{1}^{\bigstar}$, respectively. (Increasing $w^{(1,1)}$ shrinks the physical region, but leaves $w_{1}^{\prime}$ unchanged.) Notably, their scaling for $k_{T}\gg\Lambda_{\mathrm{QCD}}$ is model-independently predicted to be $w^{(N,n)}(k_{T})\sim(\Lambda_{\mathrm{QCD}}/k_{T})^{n}$ Lee et al. (2025). Therefore if heavy-quark fragmentation indeed exhibits “quantum” behavior at $k_{T}\sim\Lambda_{\mathrm{QCD}}$ as in Fig. 3, there would be the exciting possibility to observe a quantum-to-classical transition as $k_{T}$ increases.
@@ VII Conclusions and outlook
close to the isotropic limit $w_{1}=2/3$. This disagreement has, to our knowledge, not been resolved Falk (1996). We therefore strongly encourage a new measurement of $w_{1}$ using state-of-the-art experimental techniques. We also encourage experimental tests of spin correlations in nearby dihadrons and of the TMD generalizations of $w_{1}$ that we introduced, all of which possess deep physical interpretations within the new QI framework of hadronization that we have presented here.
@@ Acknowledgments:
We would like to thank Jordi Tura Brugués, Herbi Dreiner, Jonas Helsen, Piet Mulders, Marieke Postma, Iain Stewart, and Jesse Thaler for fruitful discussion, as well as the members of the Nikhef LHCb Group and Mick Mulder in particular. We gratefully acknowledge the hospitality of the Nikhef Theory Group and the Erwin Schrödinger Institute. K.L. and Z.S. were supported by the Office of Nuclear Physics of the U.S. Department of Energy under Contract No. DE-SC0011090. Z.S. was also supported by a fellowship from the MIT Department of Physics. R.v.K. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002090 COLORFREE). J.M. was supported by the D-ITP consortium, a program of NWO that is funded by the Dutch Ministry of Education, Culture and Science (OCW).
@@ VII.1.1 Factorization and general structure
as in Eq. (1), where the coefficients $\mathrm{d}\sigma_{\mathcal{O}}$ are also denoted by $\mathrm{d}\sigma$ (without subscript), $\mathrm{d}\sigma_{LL}$, and $\mathrm{d}\sigma_{TT}$, respectively. They are given in terms of dihadron fragmentation functions $D^{\mathcal{O}}_{i\to h_{1}h_{2}}$ as Metz and Vossen (2016); Pitonyak et al. (2024); Rogers et al. (2025); Pitonyak et al. (2025); Huang et al. (2025)
@@
For spin-$1/2$ baryons this results in Eq. (1) if we define the shorthands
@@
with $\tilde{D}_{TT,i}$ again defined in full analogy, where the right-hand side is evaluated at some low input scale $\mu\sim 2\,\mathrm{GeV}$. (In a more sophisticated setup, one may also define $\tilde{D}_{LL}$ and $\tilde{D}_{TT}$ as ratios of Mellin moments of the singlet and gluon or valence DiFFs to form exact renormalization-group invariants.) Eq. (2) equally well applies to $\tilde{D}_{LL,i}$ and $\tilde{D}_{TT,i}$ when defined as in Eq. (S7), where the normalized density matrix is now directly computed in terms of the renormalized fragmentation correlator for given hadron-hadron spin state at the low scale. Using Eq. (S7) in practice requires one to first determine the underlying DiFFs from a global fit to the available collider data for the yields $N_{\mathcal{O}}$ extracted from decay distributions in many different bins. Nevertheless, we believe this would be worth the effort it entails, since one would gain access to the degree of CHSH violation also as a function of e.g. the parton type.
@@ VII.1.2 Decay distribution for nearby hyperon pairs
where $\bm{P}=2\langle\hat{\bm{S}}\rangle$ with $\lvert\bm{P}\rvert\leq 1$ is the Bloch vector of the (anti)hyperon, $\bm{n_{p}}$ ($\bm{n_{\bar{p}}}$) is a unit vector pointing in the direction of the (anti)proton in the rest frame of the parent particle, $\theta_{p}$ and $\varphi_{p}$ ($\theta_{\bar{p}}$ and $\varphi_{\bar{p}}$) are the polar coordinates of $\bm{n_{p}}$ ($\bm{n_{\bar{p}}}$) in the respective rest frame, and $\alpha_{\mp}$ are the hyperon decay parameters. With the above common sign for $\alpha_{\pm}$ for hyperon and antihyperon, one has opposite signs for $\alpha_{-}=0.732\pm 0.014$ and $\alpha_{+}=-0.758\pm 0.012$ Zyla et al. (2020) (and thus $\alpha_{-}\alpha_{+}<0$), as required by the approximate CP symmetry.
@@
where the rank-two tensor $\mathsf{T}=(T_{jk})$ has components $T_{jk}/4=\langle S_{1,j}\otimes S_{2,k}\rangle$.
@@ VII.2 Inclusive fragmentation to excited heavy baryons
The factorization theorem for the production of a boosted polarized heavy hadron $H$ containing a heavy quark of mass $m_{Q}$ at strict leading power in $\Lambda_{\mathrm{QCD}}/m_{Q}$ is simply a product Mele and Nason (1991); Jaffe and Randall (1994); Falk and Peskin (1994)
@@
i.e., the normalized spin density matrix of the light constituents in any given bin of the heavy-hadron kinematics is directly given by ratios of the nonperturbative coefficients. The normalized coefficients $\chi_{H,\mathcal{O}_{\ell}}/\chi_{H}$ for a light state with angular momentum $j=1$ or $j=3/2$ are in direct correspondence to the usual Falk-Peskin parameters $w_{j}$. Specifically, for the set of operators $\{\mathcal{O}_{\ell}\}=\{\frac{\mathbb{1}}{4},\,\hat{J}_{z}^{2}\}$ allowed by the symmetries in the case $j=1$, one readily converts the coefficients to the form involving $w_{1}$ in Eq. (3) using $\hat{J}_{z}^{2}=\mathbb{1}-\lvert 0_{z}\rangle\langle 0_{z}\rvert$.
@@
Here $Y_{jm}$ is a spherical harmonic, $\theta$ ($\varphi$) is the pion polar (azimuthal) angle relative to the fragmentation axis in the candidate $\Sigma_{Q}^{(*)}\to\Lambda_{Q}\pi$ rest frame, $s=1/2$ ($3/2$) and $h$ are the total spin and spin along the $\bm{z}$ axis of the $\Sigma_{Q}$ ($\Sigma_{Q}^{*}$) baryon, $k$ is the spin of the final-state $\Lambda_{Q}$ along the $\bm{z}$ axis, and $m=h-k$. We now dress Eqs. (3) and (S13) with the appropriate Clebsch-Gordan coefficients, average over the helicities of the unpolarized initial-state heavy-quark, and sum over the helicities of the final-state $\Lambda_{Q}$ (whose polarization is not reconstructed). This results in the following angular distributions for the decay products Falk and Peskin (1994),
@@
Here we have performed the integral over the $\varphi$ dependence, which as expected is flat since no azimuthal direction is preferred. Here we assumed that the $\Sigma_{Q}$ and $\Sigma_{Q}^{*}$ form well-separated resonances $\Gamma_{\Sigma_{Q}},\Gamma_{\Sigma_{Q}^{*}}\ll m_{\Sigma_{Q}^{*}}-m_{\Sigma_{Q}}$ relative to the mass splitting in the heavy-quark spin symmetry doublet, which is reasonably well satisfied Aaltonen et al. (2012); Aaij et al. (2019). If needed, the decay distributions can readily be generalized to account for their interference Falk and Peskin (1994), and still can be expressed purely in terms of $w_{1}$ in that case. We note that the continuum $Q\to\Lambda_{Q}\pi$ background to Eq. (VII.2) is in fact described by a heavy-light DiFF that is of interest of its own, with the pion polar angle tracking its residual momentum fraction $z_{\pi}$ in this case. Importantly, however, no feature appears around $m_{\Lambda_{Q}\pi}\approx m_{\Sigma_{Q}^{*}}$ in the double-differential $(z_{\pi},m_{\Lambda_{Q}\pi})$ distribution in the background case. We therefore expect that a standard sideband subtraction will remain viable to remove the background also bin by bin in $\theta$. (Signal-background interference is likewise suppressed by the narrow widths of the resonances.) Finally, we stress that care must be taken to avoid biasing the angular distribution in Eq. (VII.2) when applying acceptance cuts on the pion transverse momentum with respect to the beam axis, as typically done in spectroscopic analyses Aaltonen et al. (2012); Aaij et al. (2019) to enrich the sample with signal events and suppress the $\Lambda_{Q}\pi$ continuum background. An attractive way to assess systematic uncertainties from correcting for the acceptance (as well as from subleading $1/m_{Q}$ corrections on the theory side) is offered by the corresponding angular distribution for $\Sigma_{Q}$ baryons, which is flat in $\cos\theta$ for unpolarized collisions at leading order in the heavy-quark expansion.

### v2

@@ VI TMD heavy-quark fragmentation
It is experimentally possible Brandenburg et al. (1997) that inclusive fragmentation $Q\to\Sigma_{Q}^{*}X$ in fact produces an isotropic spin state with $w_{1}\approx 2/3$. It is interesting to ask, therefore, whether breaking the azimuthal symmetry by measuring a transverse momentum $\bm{k}_{T}\perp\bm{z}$ provides access to richer spin-$1$ dynamics. Here the transverse momentum-dependent (TMD) observable $\bm{k}_{T}\gtrsim\Lambda_{\mathrm{QCD}}$, may e.g. be (a) the recoil of the heavy hadron pair in $e^{+}e^{-}\to\Sigma_{Q}^{*}H_{\bar{Q}}X$ von Kuk et al. (2023); Dai et al. (2024); von Kuk et al. (2024); Lee et al. (2025b), (b) the hadron transverse momentum relative to a jet containing it Kang et al. (2020); Arratia et al. (2020); Lee et al. (2025b), or (c) the transverse momentum of an additional fragmentation pion Bacchetta et al. (2023); Lee et al. (2025b). In any of these cases, the most general light angular momentum state compatible with the symmetries reads Lee et al. (2025b)
@@
where $k_{T}=\lvert\bm{k}_{T}\rvert$, $\bm{x}=\bm{k}_{T}/k_{T}$, $\bm{y}=\bm{z}\times\bm{x}$, and the $w^{(N,n)}$ are the nonperturbative coefficients. Our previous discussion elegantly carries over if we define
@@
i.e., by finding the axis $\bm{z^{\prime}}$ that maximizes the neutral polarization rate $1-w_{1}^{\prime}$. A rotation from $\bm{z}$ to $\bm{z}^{\prime}$ in fact diagonalizes $\rho$ with eigenvectors $\lvert m_{z^{\prime}}\rangle$, so we can think of $\bm{z^{\prime}}$ as a novel polarization axis that is spontaneously generated by the fragmentation dynamics. Here we can make immediate use of $\bm{z}^{\prime}$ by repeating our previous QI analysis for $w_{1}^{\prime}$, as shown in Fig. 3, where $w^{(2,0)}=0$ corresponds to $w_{1}=2/3$ after integrating over $\bm{k}_{T}$. The value of $w_{1}^{\prime}$ is driven by $w^{(2,1)}$ and $w^{(2,2)}$, and even for $w_{1}=2/3$ we find physical regions in their parameter space with nontrivial QI interpretations $w_{1}^{\prime}<1/2$ or $w_{1}^{\prime}<w_{1}^{\bigstar}$, respectively. (Increasing $w^{(1,1)}$ shrinks the physical region, but leaves $w_{1}^{\prime}$ unchanged.) Notably, their scaling for $k_{T}\gg\Lambda_{\mathrm{QCD}}$ is model-independently predicted to be $w^{(N,n)}(k_{T})\sim(\Lambda_{\mathrm{QCD}}/k_{T})^{n}$ Lee et al. (2025b). Therefore if heavy-quark fragmentation indeed exhibits “quantum” behavior at $k_{T}\sim\Lambda_{\mathrm{QCD}}$ as in Fig. 3, there would be the exciting possibility to observe a quantum-to-classical transition as $k_{T}$ increases.
@@ VII Conclusions and outlook
close to the isotropic limit $w_{1}=2/3$. This disagreement has, to our knowledge, not been resolved Falk (1996). We therefore strongly encourage a new measurement of $w_{1}$ using state-of-the-art experimental techniques. We also encourage experimental tests of spin correlations in nearby dihadrons STAR Collaboration (2025) and of the TMD generalizations of $w_{1}$ that we introduced, all of which possess deep physical interpretations within the new QI framework of hadronization that we have presented here.
@@ Acknowledgments:
We would like to thank Daniël Boer, Jordi Tura Brugués, Herbi Dreiner, Jonas Helsen, Piet Mulders, Duff Neill, Marieke Postma, Iain Stewart, and Jesse Thaler for fruitful discussion, as well as the members of the Nikhef LHCb Group and Mick Mulder in particular. The authors gratefully acknowledge the hospitality of the Nikhef Theory Group, the Institute for Theoretical Physics at the University of Amsterdam, and the Erwin Schrödinger Institute. K.L. and Z.S. were supported by the Office of Nuclear Physics of the U.S. Department of Energy under Contract No. DE-SC0011090. Z.S. was also supported by a fellowship from the MIT Department of Physics. R.v.K. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002090 COLORFREE). J.M. was supported by the D-ITP consortium, a program of NWO that is funded by the Dutch Ministry of Education, Culture and Science (OCW).
@@ VII.1.1 Factorization and general structure
as in Eq. (1), where the coefficients $\mathrm{d}\sigma_{\mathcal{O}}$ in turn are also denoted by $\mathrm{d}\sigma$ (without subscript), $\mathrm{d}\sigma_{LL}$, and $\mathrm{d}\sigma_{TT}$. They are given in terms of dihadron fragmentation functions $D^{\mathcal{O}}_{i\to h_{1}h_{2}}$ as Metz and Vossen (2016); Pitonyak et al. (2024); Rogers et al. (2025); Pitonyak et al. (2025); Huang et al. (2025)
@@
For spin-$1/2$ baryons this results in Eq. (1) if we define the shorthand
@@
with $\tilde{D}_{TT,i}$ again defined in full analogy, where the right-hand side is evaluated at some low input scale $\mu\sim 2\,\mathrm{GeV}$. (In a more sophisticated setup, one may also define $\tilde{D}_{LL}$ and $\tilde{D}_{TT}$ as ratios of Mellin moments of the singlet and gluon or valence DiFFs to form exact renormalization-group invariants.) Eq. (3) equally well applies to $\tilde{D}_{LL,i}$ and $\tilde{D}_{TT,i}$ when defined as in Eq. (S7), where the normalized density matrix is now directly computed in terms of the renormalized dihadron fragmentation correlator at the low scale. Using Eq. (S7) in practice requires one to first determine the underlying DiFFs from a global fit to the available collider data for the yields $N_{\mathcal{O}}$ extracted from decay distributions in many different bins. Nevertheless, we believe this would be worth the effort it entails, since one would gain access to the degree of CHSH violation also as a function of e.g. the parton type.
@@ VII.1.2 Decay distribution for nearby hyperon pairs
where $\bm{P}=2\langle\hat{\bm{S}}\rangle$ with $\lvert\bm{P}\rvert\leq 1$ is the Bloch vector of the (anti)hyperon, $\bm{n_{p}}$ ($\bm{n_{\bar{p}}}$) is a unit vector pointing in the direction of the (anti)proton in the rest frame of the parent particle, $\theta_{p}$ and $\varphi_{p}$ ($\theta_{\bar{p}}$ and $\varphi_{\bar{p}}$) are the polar coordinates of $\bm{n_{p}}$ ($\bm{n_{\bar{p}}}$) in the respective rest frame, and $\alpha_{\mp}$ are the hyperon decay parameters. With the above common sign in front of $\alpha_{\pm}$ for hyperon and antihyperon, one has opposite signs for $\alpha_{-}=0.732\pm 0.014$ and $\alpha_{+}=-0.758\pm 0.012$ Zyla et al. (2020) (and thus $\alpha_{-}\alpha_{+}<0$), as required by the approximate CP symmetry.
@@
where the rank-two tensor $\mathsf{T}=(T_{k\ell})=\operatorname{diag}(\tilde{D}_{TT},\tilde{D}_{TT},\tilde{D}_{LL})$ has components $T_{k\ell}/4=\langle S_{1,k}\otimes S_{2,\ell}\rangle$. We note that Eq. (VII.1.2) takes the same form for the decay products of same-sign hyperon pairs (or other exclusive baryonic decays that follow Eq. (S9)) after suitably replacing the decay constants.
Comparison to Ref. STAR Collaboration (2025):
Very recently, and after the first preprint version of this manuscript was made public, the STAR collaboration published a measurement STAR Collaboration (2025) of the differential distribution
$\frac{1}{N}\frac{\mathrm{d}N}{\mathrm{d}\cos\theta^{*}}=\frac{1}{2}\bigl{[}1+\alpha_{-}\alpha_{+}\,P_{\Lambda\bar{\Lambda}}\,\cos\theta^{*}\bigr{]}\,,$ (S11)
where in our notation $\cos\theta^{*}=\bm{n}_{p}\cdot\bm{n}_{\bar{p}}$ and the coefficient $P_{\Lambda\bar{\Lambda}}$ encodes the strength of the angular correlation between the decay (anti)protons, and thus the strength of the spin correlation between the parent (anti)hyperons. Projecting our Eq. (VII.1.2) onto the distribution differential in $\cos\theta^{*}$, we find the relation
$3P_{\Lambda\bar{\Lambda}}=\operatorname{Tr}\mathsf{T}=2\tilde{D}_{TT}+\tilde{D}_{LL}\,.$ (S12)
The measurement of $P_{\Lambda\bar{\Lambda}}$ in Ref. STAR Collaboration (2025), as well as those for same-sign hyperon pairs obtained in Ref. STAR Collaboration (2025), therefore experimentally constrain a very specific linear combination of dihadron fragmentation functions. Ref. STAR Collaboration (2025) found that for nearby dihadrons, which as discussed in the main text and the above supplemental material is the most interesting kinematic region, the spin correlation $P_{\Lambda\bar{\Lambda}}=0.181\pm 0.035_{\mathrm{stat}}\pm 0.022_{\mathrm{sys}}$ was near its maximal value of $1/3$ (with some loss expected due to feed-down contributions from higher resonances), while $P_{\Lambda\Lambda}$ and $P_{\bar{\Lambda}\bar{\Lambda}}$ were still compatible with zero. These experimental constraints are indicated in Fig. 1 as of the second preprint version of this manuscript. Notably, either of the experimental results is still compatible with entanglement or even CHSH violation in the upper left corner, since those depend on the precise breakdown between $\tilde{D}_{LL}$ and $\tilde{D}_{TT}$ and in particular the sign of $\tilde{D}_{LL}$. We stress that two caveats apply when interpreting the measurement of Ref. STAR Collaboration (2025) in this way; specifically, (a) further work is needed to ascertain the validity of the factorization in terms of fragmentation functions, which might require a more stringent cut on the $p_{T}$ of the hyperon candidates, and (b) the equivalence between the respective reference frames and angles should be carefully checked.
The analysis in Ref. STAR Collaboration (2025) was in part motivated by theory considerations that postulate a rotationally invariant distribution of decay products, as e.g. appropriate for exclusive $\Lambda\bar{\Lambda}$ production, in which case $\cos\theta^{*}$ is the only angular observable of interest. As we have discussed, however, the fragmentation process necessarily retains a dependence on the fragmentation axis $\hat{z}$, i.e., the orientation relative to the hard scattering process that produces the parent parton, and whose hard scattering products balance the parton’s color charge. For this reason the polarization tensor T need not be proportional to the identity, and indeed in general has two nontrivial degrees of freedom ($\tilde{D}_{LL}$ and $\tilde{D}_{TT}$) according to our symmetry analysis that led to Eq. (1). We therefore strongly encourage an extension of the exciting measurement reported in Ref. STAR Collaboration (2025) to the more differential angular distribution in Eq. (VII.1.2), which would break the degeneracy between $\tilde{D}_{LL}$ and $\tilde{D}_{TT}$ and hold the potential of a definitive interpretation in terms of QI constraints on possible hadronization models. Alternatively, the degeneracy could at this point also be broken by a simpler, direct measurement of $\tilde{D}_{LL}$ via the $\cos\theta_{p}\cos\theta_{\bar{p}}$ correlation in Eq. (VII.1.2) in a way that is inclusive over $\phi_{p,\bar{p}}$, i.e.,
$\frac{1}{N}\frac{\mathrm{d}N}{\mathrm{d}\cos\theta_{p}\,\mathrm{d}\cos\theta_{\bar{p}}}$ $=\frac{1}{4}\bigl{[}1+\alpha_{-}\alpha_{+}\,\tilde{D}_{LL}\,\cos\theta_{p}\,\cos\theta_{\bar{p}}\bigr{]}\,.$ (S13)
@@ VII.2 Inclusive fragmentation to excited heavy baryons
The factorization theorem for the production of a boosted polarized heavy hadron $H$ containing a heavy quark of mass $m_{Q}$ at strict leading power in $\Lambda_{\mathrm{QCD}}/m_{Q}$ is simply a product Mele and Nason (1991); Jaffe and Randall (1994); Falk and Peskin (1994),
@@
i.e., the normalized spin density matrix of the light constituents in any given bin of the heavy-hadron kinematics is directly given by ratios of the nonperturbative coefficients. The normalized coefficients $\chi_{H,\mathcal{O}_{\ell}}/\chi_{H}$ for a light state with angular momentum $j=1$ or $j=3/2$ are in direct correspondence to the usual Falk-Peskin parameters $w_{j}$. Specifically, for the set of operators $\{\mathcal{O}_{\ell}\}=\{\mathbb{1}/4,\,\hat{J}_{z}^{2}\}$ allowed by the symmetries in the case $j=1$, one readily converts the coefficients to the form involving $w_{1}$ in Eq. (4) using $\hat{J}_{z}^{2}=\mathbb{1}-\lvert 0_{z}\rangle\langle 0_{z}\rvert$.
@@
Here $Y_{jm}$ is a spherical harmonic, $\theta$ ($\varphi$) is the pion polar (azimuthal) angle relative to the fragmentation axis in the candidate $\Sigma_{Q}^{(*)}\to\Lambda_{Q}\pi$ rest frame, $s=1/2$ ($3/2$) and $h$ are the total spin and spin along the $\bm{z}$ axis of the $\Sigma_{Q}$ ($\Sigma_{Q}^{*}$) baryon, $k$ is the spin of the final-state $\Lambda_{Q}$ along the $\bm{z}$ axis, and $m=h-k$. We recall the following, very simple form of the fragmentation axis $\bm{z}$ when expressed in terms of the beam three-momenta $\bm{P}_{a,b}$ in the candidate rest frame,
$\bm{z}=-\frac{\bm{P_{a}}+\bm{P_{b}}}{\lvert\bm{P_{a}}+\bm{P_{b}}\rvert}\,,$ (S17)
where the beams can be protons or leptons. We now dress Eqs. (4) and (S16) with the appropriate Clebsch-Gordan coefficients, average over the helicities of the unpolarized initial-state heavy-quark, and sum over the helicities of the final-state $\Lambda_{Q}$ (whose polarization is not reconstructed). This results in the following angular distributions for the decay products Falk and Peskin (1994),
@@
Here we have performed the integral over the $\varphi$ dependence, which as expected is flat since no azimuthal direction is preferred. We assumed that the $\Sigma_{Q}$ and $\Sigma_{Q}^{*}$ form well-separated resonances $\Gamma_{\Sigma_{Q}},\Gamma_{\Sigma_{Q}^{*}}\ll m_{\Sigma_{Q}^{*}}-m_{\Sigma_{Q}}$ relative to the mass splitting in the heavy-quark spin symmetry doublet, which is reasonably well satisfied Aaltonen et al. (2012); Aaij et al. (2019). If needed, the decay distributions can readily be generalized to account for their interference Falk and Peskin (1994), and still can be expressed purely in terms of $w_{1}$ in that case. We note that the continuum $Q\to\Lambda_{Q}\pi$ background to Eq. (VII.2) initiated by a boosted heavy quark is in fact described by a heavy-light DiFF that is of interest of its own, and upon expansion in $\Lambda_{\mathrm{QCD}}/m_{Q}\ll 1$ takes the same form as Eq. (S14) with a more differential nonperturbative matrix element $\chi_{\Lambda_{Q}\pi}$,
$\frac{\mathrm{d}\sigma_{\Lambda_{Q}\pi X}^{\mathcal{O}_{\ell}}}{\mathrm{d}^{3}\bm{P}_{\Lambda_{Q}\pi}^{\mathrm{lab}}\,\mathrm{d}m_{\Lambda_{Q}\pi}\,\mathrm{d}\cos\theta}=\frac{\mathrm{d}\sigma_{QX}}{\mathrm{d}^{3}\bm{P}_{Q}^{\mathrm{lab}}}\,\int\!\mathrm{d}E_{\pi}\,\mathrm{d}p_{\pi}^{z}\,\chi_{\Lambda_{Q}\pi}(E_{\pi},p_{\pi}^{z})\,\delta\bigl{(}m_{\Lambda_{Q}\pi}-m_{\Lambda_{Q}}-m_{\pi}-E_{\pi}\bigr{)}\,\delta\Bigl{(}\cos\theta-\frac{p_{\pi}^{z}}{\lvert\bm{p}_{\pi}\rvert}\Bigr{)}\,,$ (S19)
where $m_{\Lambda_{Q}\pi}$ is the invariant mass of the candidate pair and $E_{\pi}$ and $\bm{p}_{\pi}$ are the energy and three-momentum of the pion in the candidate center-of-mass frame, $\bm{p}_{\pi}=-\bm{p}_{\Lambda_{Q}}$. (For clarity, we have made explicit that the approximately identical kinematics $\bm{P}_{\Lambda_{Q}\pi}^{\mathrm{lab}}=\bm{P}_{Q}^{\mathrm{lab}}+\mathcal{O}(\Lambda_{\mathrm{QCD}})$ of the candidate pair and the heavy-quark progenitor are evaluated in the lab frame instead; in practice this dependence will always be integrated over a certain bin at the level of the prefactor.) The dependence of $\chi_{\Lambda_{Q}\pi}$ on $p_{\pi}^{z}$, which determines the $\cos\theta$ shape of the background, and the correlation between $p_{\pi}^{z}$ and $E_{\pi}$, satisfy the obvious on-shell constraints $E_{\pi}\geq m_{\pi}$ and $\lvert p_{\pi}^{z}\rvert\leq\sqrt{E_{\pi}^{2}-m_{\pi}^{2}}$, but are complicated in general. Importantly, however, nothing privileges the point $E_{\pi}\approx m_{\Sigma_{Q}^{*}}-m_{\Lambda Q}-m_{\pi}$ in the double-differential continuum fragmentation function $\chi_{\Lambda_{Q}\pi}$, and thus no feature appears around $m_{\Lambda_{Q}\pi}\approx m_{\Sigma_{Q}^{*}}$ in the double-differential $(m_{\Lambda_{Q}\pi},\cos\theta)$ distribution in the background case. We therefore expect that a standard sideband subtraction will remain viable to remove the background also bin by bin in $\theta$. (Signal-background interference is likewise suppressed by the narrow widths of the resonances.) Finally, we stress that care must be taken to avoid biasing the angular distribution in Eq. (VII.2) when applying acceptance cuts on the pion transverse momentum with respect to the beam axis, as typically done in spectroscopic analyses Aaltonen et al. (2012); Aaij et al. (2019) to enrich the sample with signal events and suppress the $\Lambda_{Q}\pi$ continuum background. An attractive way to assess systematic uncertainties from correcting for the acceptance (as well as from subleading $1/m_{Q}$ corrections on the theory side) is offered by the corresponding angular distribution for $\Sigma_{Q}$ baryons, which is flat in $\cos\theta$ for unpolarized collisions at leading order in the heavy-quark expansion.